# Groupoid Graded Semisimple Rings

### v1

@@ 1. Introduction
Semisimple rings and the Wedderburn-Artin Theorem lie at the basis of classical Ring Theory. Later, the study of group graded structures became important for the development of different branches of Mathematics [31], [17], [11]. Many classical concepts and results in Ring and Module Theory have a group graded version, for example the Wedderburn-Artin Theorem [24], [31, Theorem 2.10.10] or [10, Theorem 2.12]. Nowadays, groupoids [18] play an important role in Mathematics [33], [28], [12], and groupoid graded structures have already appeared in different contexts such as partial actions [2], [30] and Leavitt path algebras [16].
A groupoid, together with a zero, can be regarded as a semigroup. Thus, some results about group graded rings might be shrouded in the theory of semigroup graded rings. The first work that begun the study of groupoid graded rings and modules as a separate class was [26]. Usually, first results in the theory were obtained supposing that either the graded ring has an identity element and/or that the groupoid has a finite number of objects [25], [27], [34]. A systematic study of groupoid graded rings and modules that do not satisfy these conditions begun in [8] where also the first results about groupoid graded semisimple modules were given. In that work, graded rings are supposed to have a good set of local units. They refer to such rings as object unital rings, a term that we adopt in this work (see Subsection 2.3). This concept was already used in [32] under the name of locally unital ring.
@@
Naturally, in the groupoid graded context, rings do not necessarily have a unity. In fact, imposing the existence of an identity element implies that there is only a finite number of idempotents of the groupoid inside the support of the graded ring [26, Proposition 2.1.1]. The possibility of an infinite number of idempotents in a groupoid makes that some classical definitions do not generalize as directly as they do in the group graded case. For example, many interesting examples of groupoid graded rings are decomposed as an infinite direct sum of nonzero graded right modules. The usual definition of artinianity does not apply in this case, but another descending chain condition is useful. We will talk about $\Gamma_{0}$-artinian rings, see Subsection 5.2. In other cases, the natural generalization of certain concepts lead to unexpected properties. For example, groupoid graded division rings are not necessarily graded prime rings. Other definitions are not useful in the groupoid graded context. For example, if a module is graded by a groupoid but not by a group, then no element is linearly independent. For this, we define pseudo-free modules instead of free modules, see Subsection 3.1 for more details.
Other difficulty that appears when working with groupoid graded ring theory is how to define groupoid gradings on matrix rings. In the group graded context, we need a finite sequence of elements of a group to obtain a group grading in a matrix ring (of finite size). In our case, we will need sequences of suitable subsets of the groupoid in order to obtain groupoid gradings in matrix rings. On the other hand, these graded matrix rings have similar properties to those of the group graded context.
@@
It is a well-known result in Ring Theory that every unital right (left) module over a division ring has a basis (see, for example, [19, Theorem 2.4]). One can use the Wedderburn-Artin theorem to prove the converse. In fact, suppose that $R$ is a ring with unity such that every unital right (left) $R$-module is free. This implies that every unital right (left) $R$-module is projective. By [21, Theorem 2.8], $R$ is a semisimple ring and it follows that $R$ must be a finite product of matrix rings over division rings. But all modules over such ring are free if and only if the ring is a division ring. Analogous arguments show the correspondent characterization of group graded division rings (cf. [4, Theorem 3.3] for one implication). We were surprised to build a non sophisticated example of a groupoid graded ring (a $3\times 3$ matrix ring with a suitable grading) over which all graded right modules are pseudo-free, but that is not a graded division ring. With this example in mind and using our graded Wedderburn-Artin Theorem, we are able to characterize the graded rings all of whose modules are pseudo-free.
@@
Section 3 is dedicated to study groupoid gradings on matrix rings, homomorphisms between groupoid graded modules and the identification of graded matrix rings with graded endomorphism rings of certain graded modules. We begin this section talking about pseudo-free modules. Next, we define and prove some results about gr-homomorphisms and homomorphisms with degree. Then, we show how to obtain groupoid gradings in matrix rings over graded rings. One of the main results of this section is the expression of homomorphisms with degree as matrices. Later, in order to obtain results about left modules from those about right modules, we consider the opposite ring of a graded ring as a graded ring. These techniques are useful to get isomorphisms between categories of graded left modules over a graded ring and graded right modules over the opposite ring. We also show other isomorphisms between categories of graded modules that are obtained when the groupoid is connected. We finish this third section with groupoid gradings in additive groups of rectangular matrices.
@@
Example 2.3 tells us how to obtain a groupoid graded ring from a small preadditive category. This example and [29, 15] motivate us to apply the concepts discussed in previous sections to the categorical context. This is the main objective of Section 8. We begin showing that additive contravariant functors from a small preaddtive category to abelian groups can be regarded as graded modules over the ring of the category. After that, we obtain new characterizations of semisimple categories. We also define and characterize simple artinian categories and division categories. We show that the concept of free functors given in [29] is linked with our concept of pseudo-free modules and we use this to get a characterization of categories for which all functors to abelian groups are free.
@@ 2.1. Groupoids
    Let $X$ be a nonempty set. We endow $\Gamma:=X\times X$ with a structure of groupoid in the following way. For $(y,x)\in\Gamma$, define $d(y,x)=x$, $r(y,x)=y$ and $(y,x)^{-1}=(x,y)$. Thus $\Gamma_{0}=X$ and there exists a unique morphism between two objects $x,y\in\Gamma_{0}$. For each $(y,x),(z,w)\in\Gamma$, the composition $(z,w)(y,x)$ is defined if, and only if, $w=y$, and in this event, $(z,y)(y,x)=(z,x)$.
@@
    Generalizing the previous example, let $X$ be a nonempty set and $G$ be a group. One can endow $\Gamma:=X\times G\times X$ with a structure of groupoid in the following way. For each $(y,g,x)\in\Gamma$, define $d(y,g,x)=x$, $r(y,g,x)=y$ and $(y,g,x)^{-1}=(x,g^{-1},y)$. Thus $\Gamma_{0}=X$ and the morphisms between two objects $x,y\in\Gamma_{0}$ can be identified with the elements of $G$. For each $(z,h,w),(y,g,x)\in\Gamma$, the composition $(z,h,w)(y,g,x)$ is defined if, and only if, $w=y$, and in this event, $(z,h,y)(y,g,x)=(z,hg,x)$. Notice that we obtain the foregoing example when $G$ is the trivial group.
    This example is important because any groupoid is the disjoint union of subgroupoids isomorphic (in a non canonical way) to the groupoid presented in this example, see for example [6, p. 125].∎
@@ 2.2. Groupoid graded additive groups
We continue fixing some general notation. Let $X=\bigoplus_{\gamma\in\Gamma}X_{\gamma}$ be a $\Gamma$-graded additive group. For each $\gamma\in\Gamma$, $X_{\gamma}$ is called the homogeneous component of degree $\gamma$ of $X$, its nonzero elements are said to be homogeneous of degree $\gamma$ and we write $\deg(x)=\gamma\in\Gamma$ when $0\neq x\in X_{\gamma}$. For each $\gamma\in\Gamma$ and $x\in X$, we write $x=\sum_{\gamma\in\Gamma}x_{\gamma}$ with $x_{\gamma}\in X_{\gamma}$ (with all but finitely many $x_{\gamma}$ nonzero) and we call $x_{\gamma}$ the homogeneous component of degree $\gamma$ of $x$. The set of the homogeneous elements of $X$ is $h(X):=\bigcup_{\gamma\in\Gamma}X_{\gamma}$. We define the support of $X$ as $\operatorname{supp}(X):=\{\gamma\in\Gamma:X_{\gamma}\neq 0\}$ and the support of $x\in X$ as $\operatorname{supp}(x):=\{\gamma\in\Gamma:x_{\gamma}\neq 0\}$. When $\sigma,\tau\in\Gamma$ are such that $d(\sigma)\neq r(\tau)$, that is, $\sigma\tau$ is not defined, we will adopt the convention $X_{\sigma\tau}:=\{0\}$. If $Y$ is a $\Gamma$-graded additive group, then a homomorphism of groups $g:X\to Y$ it is said to be a gr-homomorphism of groups if $g(X_{\sigma})\subseteq Y_{\sigma}$ for all $\sigma\in\Gamma$. If, moreover, $g$ is bijective, then we say that $g$ is a gr-isomorphism of groups, that $X$ is gr-isomorphic to $Y$ as groups and we will denote it by $X\cong_{gr}Y$.
@@ 2.3. Groupoid graded rings
Let $R$ be a ring. We say that $R$ is a $\Gamma$-graded ring if there is a family $\{R_{\gamma}\}_{\gamma\in\Gamma}$ of additive subgroups of $R$ such that $R=\bigoplus\limits_{\gamma\in\Gamma}R_{\gamma}$ and $R_{\gamma}R_{\delta}\subseteq R_{\gamma\delta}$, for each $\gamma,\delta\in\Gamma$. Following [7], we say that $R$ is object unital if, for all $e\in\Gamma_{0}$, the ring $R_{e}$ is unital with identity element $1_{e}$, and for all $\gamma\in\Gamma$ and $r\in R_{\gamma}$, we have $1_{r(\gamma)}r=r1_{d(\gamma)}=r$. This concept has received other names in the semigroup graded context, see [1]. It follows from [26, Proposition 2.1.1] that the object unital $\Gamma$-graded ring $R=\bigoplus_{\gamma\in\Gamma}R_{\gamma}$ is unital if, and only if, $\Gamma_{0}^{\prime}=\{e\in\Gamma_{0}\colon 1_{e}\neq 0\}$ is finite. In this event, $1_{R}=\sum_{e\in\Gamma_{0}^{\prime}}1_{e}$.
@@
    If $R$ is a group graded (unital) ring, then $R$ is a (object unital) groupoid graded ring. More generally, if $I$ is a nonempty set and, for each $i\in I$, $R_{i}$ is a (unital) ring graded by the group $G_{i}$, then $R:=\bigoplus_{i\in I}R_{i}$ is (object unital) graded by the groupoid obtained as the disjoint union of the groups $G_{i}$, $i\in I$, via $R_{g_{i}}:=(R_{i})_{g_{i}}$ if $g_{i}\in G_{i}$.
@@
    By [7, Proposition 16], $A\rtimes^{\alpha}_{\beta}\Gamma$ is an (object unital) groupoid graded ring via $(A\rtimes^{\alpha}_{\beta}\Gamma)_{\sigma}:=A_{r(\sigma)}u_{\sigma}$.
@@
    In [2, Section 3], given a partial action $\alpha=(\{D_{\gamma}\}_{\gamma}\in\Gamma,\{\alpha_{\gamma}\}_{\gamma\in\Gamma})$ of an ordered groupoid $\Gamma$ on a ring $R$, the skew groupoid ring $R\ast_{\alpha}\Gamma$ is defined. It is $\Gamma$-graded and it is shown to be associative and object unital under mild conditions.
@@
    In [16], it is presented a free groupoid grading on Leavitt path algebras. ∎
Now we provide what we believe is the most important class o examples of groupoid graded rings. As far as we know, this family of examples has not been studied as graded rings until now. The first example will be very important in Section 8.
@@
    Let $\mathcal{C}$ be a small preadditive category, i.e., $Obj(\mathcal{C})$ is a set and every morphism set has an additive group structure such that the composition of morphisms is $\mathbb{Z}$-bilinear. We define the ring of the category $\mathcal{C}$ as  [^1] Our definition of preadditive category is called additive category in [29, p. 9]. Our definition of ring of a category is the Gabriel functor ring of [13, p. 123] and is denoted by $\mathbf{Z}[\mathcal{C}]$ in [15, p. 346]. We point out that our ring of a category is not the same as the one in [29, Section 7].
@@
    where, given morphisms $f$ and $g$, the product $fg$ is defined by $f\circ g$ if the composition is possible and 00 otherwise. The ring (possibly without unity) $R[\mathcal{C}]$ has a natural structure of ring graded by the groupoid $Obj(\mathcal{C})\times Obj(\mathcal{C})$ via
@@ 2.4. Groupoid graded modules
A submodule $N$ of $M$ is a graded submodule if $N=\bigoplus_{\gamma\in\Gamma}N_{\gamma}$ where $N_{\gamma}:=N\cap M_{\gamma}$ for each $\gamma\in\Gamma$. Equivalently, if $n_{\gamma_{1}}+\cdots+n_{\gamma_{t}}\in N$ with $n_{\gamma_{i}}\in M_{\gamma_{i}}$ and $\gamma_{i}\neq\gamma_{j}$ for different $i,j=1,...,t$, then $n_{\gamma_{i}}\in N$ for all $i=1,...,t$.
@@
Let $M$, $N$ be $\Gamma$-graded right $R$-modules. A homomorphism of modules $g:M\to N$ it is said to be a gr-homomorphism of modules if $g(M_{\sigma})\subseteq N_{\sigma}$ for all $\sigma\in\Gamma$. We denote by $\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ the abelian group consisting of the gr-homomorphisms of modules $g:M\to N$. We also define the ring $\operatorname{End_{gr-\mathit{R}}}(M):=\operatorname{Hom_{gr-\mathit{R}}}(M,M)$. It is easy to show that if $g\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$, then $\ker g$ is a graded submodule of $M$ and $\operatorname{im}g$ is a graded submodule of $N$. A bijective gr-homomorphism of modules $g:M\to N$ is called a gr-isomorphism of modules. We say that $M$ is gr-isomorphic to $N$ as modules (denoted by $M\cong_{gr}N$) if there exists a gr-isomorphism of modules $g:M\to N$.
@@
[8, Proposition 5] Let $R$ be a $\Gamma$-graded ring and let $M$ be a $\Gamma$-graded right $R$-module. Then $M$ is unital if and only if the equality $m_{\sigma}1_{d(\sigma)}=m_{\sigma}$ holds for all $\sigma\in\Gamma$ and $m_{\sigma}\in M_{\sigma}$. ∎
In this work, unless otherwise stated (Section 6), all groupoid graded modules are supposed to be unital. Moreover, in expressions like “let $M$ be a $\Gamma$-graded $R$-module” we are assuming that $M$ is a $\Gamma$-graded right $R$-module.
@@
    Let $I$ be a set and $M$ be an $I\times I$-graded $R$-module. For each $i_{0}\in I$, every unital right $R$-module $M$ can be regarded as an $I\times I$-graded $R$-module via $M_{(i_{0},i)}:=M1_{(i,i)}$ for each $i\in I$. Furthermore, any unital $(R,R)$-bimodule (i.e. $R\cdot M=M=M\cdot R$) can be regarded as an $I\times I$-graded $(R,R)$-bimodule via $M_{(i,j)}:=1_{(i,i)}M1_{(j,j)}$, for each $i,j\in I$. In particular this applies to the ring $R[\mathcal{C}]$ of a small preadditive category $\mathcal{C}$. In this case, every ideal of $R[\mathcal{C}]$ is $\mathcal{C}_{0}\times\mathcal{C}_{0}$-graded.
@@
Let $M$ be a $\Gamma$-graded right $R$-module and $\sigma\in\Gamma$. For each $\gamma\in\Gamma$, define $M(\sigma)_{\gamma}:=M_{\sigma\gamma}$. The shift of $M$ by $\sigma$ is the $\Gamma$-graded right $R$-module $M(\sigma):=\bigoplus_{\gamma\in\Gamma}M(\sigma)_{\gamma}$. Note that $\operatorname{supp}(M(\sigma))\subseteq d(\sigma)\Gamma$. Analogously, if $M$ is a $\Gamma$-graded left $R$-module and $\gamma\in\Gamma$, set $((\sigma)M)_{\gamma}:=M_{\gamma\sigma}$. The shift of $M$ by $\sigma$ is the $\Gamma$-graded left $R$-module $(\sigma)M:=\bigoplus_{\gamma\in\Gamma}((\sigma)M)_{\gamma}$. Notice that $\operatorname{supp}((\sigma)M)\subseteq\Gamma r(\sigma)$. The following property of shifts of modules will be very useful. For a left version, see [8, Proposition 10(a)].
@@
Let $M$ be a $\Gamma$-graded $R$-module and $\sigma\in\Gamma$. Then $M(\sigma)$ equals $M(r(\sigma))$ as $R$-modules. More precisely, $M(\sigma)$ and $M(r(\sigma))$ have the same homogeneous components (although labelled with a different degree).
@@
Note that if $R$ is a $\Gamma$-graded ring and $M$ is a $\Gamma$-graded right $R$-module, then $M=\bigoplus_{e\in\Gamma_{0}}M(e)$ This decomposition will be important, so we define
@@
The following results follow immediately from the definition (see [8, Proposition 13(c)]).

### v2

@@ 1. Introduction
Semisimple rings and the Wedderburn-Artin Theorem lie at the basis of classical Ring Theory. Later, the study of group graded structures became important for the development of different branches of Mathematics [34], [20], [14]. Many classical concepts and results in Ring and Module Theory have a group graded version, for example the Wedderburn-Artin Theorem [27], [34, Theorem 2.10.10] or [13, Theorem 2.12]. Nowadays, groupoids [21] play an important role in Mathematics [36], [31], [15], and groupoid graded structures have already appeared in different contexts such as partial actions [4], [5], [33] and Leavitt path algebras [19].
A groupoid, together with a zero, can be regarded as a semigroup. Thus, some results about group graded rings might be shrouded in the theory of semigroup graded rings. To knowledge of the authors, the first work that began the study of groupoid graded rings and modules as a separate class was [29]. Usually, first results in the theory were obtained supposing that either the graded ring has an identity element and/or that the groupoid has a finite number of objects [28], [30], [37]. A systematic study of groupoid graded rings and modules that do not satisfy these conditions began in [11] where the first results about groupoid graded semisimple modules were also given. In that work, graded rings are supposed to have a good set of local units. They refer to such rings as object unital rings, a term that we adopt in this work (see Subsection 2.3). This concept was already used in [35] under the name of locally unital ring.
@@
Naturally, in the groupoid graded context, rings do not necessarily have a unity. In fact, imposing the existence of an identity element implies that there is only a finite number of idempotents of the groupoid inside the support of the graded ring [29, Proposition 2.1.1]. The possibility of an infinite number of idempotents in a groupoid does not allow some classical definitions to be generalized as directly as they do in the group graded case. For instance, many interesting examples of groupoid graded rings are decomposed as an infinite direct sum of nonzero graded right modules. The usual definition of artinianity does not apply in this case, but another descending chain condition is useful. It is the natural generalization of the notion of a categorically artinian ring, introduced in [2, Definition 1.1] and [1, Section 4.2], to the setting of groupoid graded rings. We will talk about $\Gamma_{0}$-artinian rings, see Subsection 5.2. In other cases, the natural generalization of certain concepts lead to unexpected properties. For example, groupoid graded division rings are not necessarily graded prime rings. Other definitions are not useful in the groupoid graded context. Also, if a module is graded by a groupoid but not by a group, then no element is linearly independent. For this reason, we define pseudo-free modules instead of free modules, see Subsection 3.1 for more details.
Another difficulty that appears when dealing with rings $R$ graded by a groupoid $\Gamma$ is how to induce $\Gamma$-gradings on matrix rings with entries in $R$. In the group graded context, a finite sequence of $n$ elements of the group is enough to produce a group grading on the matrix ring (of finite size $n\times n$). In our case, we will require sequences of suitable subsets of the groupoid in order to obtain groupoid gradings on matrix rings with entries in a groupoid graded ring. On the other hand, these graded matrix rings have similar properties to those that appear in the group graded context.
@@
It is a well-known result in Ring Theory that every unital right (left) module over a division ring has a basis (see, for example, [22, Theorem 2.4]). One can use the Wedderburn-Artin theorem to prove the converse. In fact, suppose that $R$ is a ring with unity such that every unital right (left) $R$-module is free. This implies that every unital right (left) $R$-module is projective. By [24, Theorem 2.8], $R$ is a semisimple ring and it follows that $R$ must be a finite product of matrix rings over division rings. But all modules over such ring are free if and only if the ring is a division ring. Analogous arguments show the correspondent characterization of group graded division rings (cf. [7, Theorem 3.3] for one implication). We were surprised to build a non-sophisticated example of a groupoid graded ring (a $3\times 3$ matrix ring with a suitable grading) over which all graded right modules are pseudo-free, but that is not a graded division ring. With this example in mind and using our graded Wedderburn-Artin Theorem, we are able to characterize graded rings whose modules are all pseudo-free.
@@
Section 3 is dedicated to study groupoid gradings on matrix rings, homomorphisms between groupoid graded modules and the identification of graded matrix rings with graded endomorphism rings of certain graded modules. We begin this section by presenting pseudo-free modules. Next, we define and prove some results about gr-homomorphisms and homomorphisms with degree. Then, we describe a key method to obtain groupoid gradings in matrix rings over graded rings. One of the main results of this section is the expression of homomorphisms with degree as matrices. Later, in order to obtain results about left modules from those about right modules, we consider the opposite ring of a graded ring as a graded ring. These techniques are useful to get isomorphisms between categories of graded left modules over a graded ring and graded right modules over the opposite ring. We also show other isomorphisms between categories of graded modules that are obtained when the groupoid is connected. We finish this third section with groupoid gradings in additive groups of rectangular matrices.
@@
Example 2.3 tells us how to obtain a groupoid graded ring from a small preadditive category. This example and [32, 18] motivate us to apply the concepts discussed in previous sections to the categorical context. This is the main objective of Section 8. We begin by showing that additive contravariant functors from a small preaddtive category to abelian groups can be regarded as graded modules over the ring of the category. After that, we obtain new characterizations of semisimple categories. We also define and characterize simple artinian categories and division categories. We show that the concept of free functors given in [32] is linked with our concept of pseudo-free modules and we use this to get a characterization of categories for which all functors to abelian groups are free.
@@ 2.1. Groupoids
    Let $X$ be a non-empty set. We endow $\Gamma:=X\times X$ with a structure of groupoid in the following way. For $(y,x)\in\Gamma$, define $d(y,x)=(x,x)$, $r(y,x)=(y,y)$ and $(y,x)^{-1}=(x,y)$. Thus $\Gamma_{0}=\{(x,x)\colon x\in X\}$ and it can be identified with $X$ in the natural way. Note that there exists a unique morphism between two objects in $\Gamma_{0}$. For each $(y,x),(z,w)\in\Gamma$, the composition $(z,w)(y,x)$ is defined if and only if $w=y$. In that case, $(z,y)(y,x)=(z,x)$.
@@
    Generalizing the previous example, let $X$ be a non-empty set and $G$ be a group with identity element $e$. We endow $\Gamma:=X\times G\times X$ with a structure of groupoid in the following way. For each $(y,g,x)\in\Gamma$, define $d(y,g,x)=(x,e,x)$, $r(y,g,x)=(y,e,y)$ and $(y,g,x)^{-1}=(x,g^{-1},y)$. Thus $\Gamma_{0}=\{(x,e,x)\colon x\in X\}$ and it can be identified with $X$. Note that the morphisms between two objects in $\Gamma_{0}$ can be identified with the elements of $G$. For each $(z,h,w),(y,g,x)\in\Gamma$, the composition $(z,h,w)(y,g,x)$ is defined if and only if $w=y$. In that case, $(z,h,y)(y,g,x)=(z,hg,x)$. Notice that we obtain the foregoing example when $G$ is the trivial group.
    This example is important because any groupoid is the disjoint union of subgroupoids isomorphic (in a non-canonical way) to the groupoid presented in this example, see for example [9, p. 125].∎
@@
Note that if $e\in\Gamma_{0}$, then $e\Gamma e=\{\gamma\in\Gamma\colon d(\gamma)=r(\gamma)=e\}$ is a group with identity element $e$. While $\Gamma(e)$ is the standard notation for this group in the literature, we choose the more suggestive notation $e\Gamma e$ to better serve our purposes and to avoid potential confusion with other concepts.
@@ 2.2. Groupoid graded additive groups
We continue fixing some general notation. Let $X=\bigoplus_{\gamma\in\Gamma}X_{\gamma}$ be a $\Gamma$-graded additive group. For each $\gamma\in\Gamma$, $X_{\gamma}$ is called the homogeneous component of degree $\gamma$ of $X$, its nonzero elements are said to be homogeneous of degree $\gamma$ and we write $\deg(x)=\gamma\in\Gamma$ when $0\neq x\in X_{\gamma}$. For each $\gamma\in\Gamma$ and $x\in X$, we write $x=\sum_{\gamma\in\Gamma}x_{\gamma}$ with $x_{\gamma}\in X_{\gamma}$ (with all but finitely many $x_{\gamma}$ nonzero) and we call $x_{\gamma}$ the homogeneous component of degree $\gamma$ of $x$. The set of the homogeneous elements of $X$ is $\operatorname{h}(X):=\bigcup_{\gamma\in\Gamma}X_{\gamma}$. We define the support of $X$ as $\operatorname{supp}(X):=\{\gamma\in\Gamma:X_{\gamma}\neq 0\}$ and the support of $x\in X$ as $\operatorname{supp}(x):=\{\gamma\in\Gamma:x_{\gamma}\neq 0\}$. When $\sigma,\tau\in\Gamma$ are such that $d(\sigma)\neq r(\tau)$, that is, $\sigma\tau$ is not defined, we adopt the convention $X_{\sigma\tau}:=\{0\}$. If $Y$ is a $\Gamma$-graded additive group, then a homomorphism of groups $g:X\to Y$ it is said to be a gr-homomorphism of groups if $g(X_{\sigma})\subseteq Y_{\sigma}$ for all $\sigma\in\Gamma$. If, moreover, $g$ is bijective, then we say that $g$ is a gr-isomorphism of groups, that $X$ is gr-isomorphic to $Y$ as groups and we will denote it by $X\cong_{gr}Y$.
@@ 2.3. Groupoid graded rings
Let $R$ be a ring. We say that $R$ is a $\Gamma$-graded ring if there is a family $\{R_{\gamma}\}_{\gamma\in\Gamma}$ of additive subgroups of $R$ such that $R=\bigoplus\limits_{\gamma\in\Gamma}R_{\gamma}$ and $R_{\gamma}R_{\delta}\subseteq R_{\gamma\delta}$, for each $\gamma,\delta\in\Gamma$. Following [10], we say that $R$ is object unital if, for all $e\in\Gamma_{0}$, the ring $R_{e}$ is unital with identity element $1_{e}$, and for all $\gamma\in\Gamma$ and $r\in R_{\gamma}$, we have $1_{r(\gamma)}r=r1_{d(\gamma)}=r$. This concept has received other names in the semigroup graded context, see [3]. It follows from [29, Proposition 2.1.1] that the object unital $\Gamma$-graded ring $R=\bigoplus_{\gamma\in\Gamma}R_{\gamma}$ is unital if and only if $\Gamma_{0}^{\prime}=\{e\in\Gamma_{0}\colon 1_{e}\neq 0\}$ is finite. In this event, $1_{R}=\sum_{e\in\Gamma_{0}^{\prime}}1_{e}$.
@@
    If $R$ is a group graded (unital) ring, then $R$ is a (object unital) groupoid graded ring. More generally, if $I$ is a non-empty set and, for each $i\in I$, $R_{i}$ is a (unital) ring graded by the group $G_{i}$, then $R:=\bigoplus_{i\in I}R_{i}$ is (object unital) graded by the groupoid obtained as the disjoint union of the groups $G_{i}$, $i\in I$, via $R_{g_{i}}:=(R_{i})_{g_{i}}$ if $g_{i}\in G_{i}$.
@@
    By [10, Proposition 16], $A\rtimes^{\alpha}_{\beta}\Gamma$ is an (object unital) groupoid graded ring via $(A\rtimes^{\alpha}_{\beta}\Gamma)_{\sigma}:=A_{r(\sigma)}u_{\sigma}$.
@@
    Given a partial action $\alpha=(\{D_{\gamma}\}_{\gamma}\in\Gamma,\{\alpha_{\gamma}\}_{\gamma\in\Gamma})$ of a groupoid $\Gamma$ on a ring $R$ as in [5, Section 3], the skew groupoid ring $R\ast_{\alpha}\Gamma$ is defined. It is $\Gamma$-graded, associative and object unital. See also [4, Section 3] for mild conditions under which $R\ast_{\alpha}\Gamma$ is associative and object unital.
@@
    In [19], a free groupoid grading on Leavitt path algebras is presented. ∎
Now we provide what we believe is the most important class of examples of groupoid graded rings. As far as we know, this family of examples has not been studied as graded rings until now. The first example will be very important in Section 8.
@@
    Let $\mathcal{C}$ be a small preadditive category, i.e., $Obj(\mathcal{C})$ is a set and every morphism set has an additive group structure such that the composition of morphisms is $\mathbb{Z}$-bilinear. We define the ring of the category $\mathcal{C}$ as  [^1] Our definition of preadditive category is called additive category in [32, p. 9]. Our definition of ring of a category is the Gabriel functor ring of [16, p. 123] and is denoted by $\mathbf{Z}[\mathcal{C}]$ in [18, p. 346]. We point out that our ring of a category is not the same as the one in [32, Section 7].
@@
    where, given morphisms $f$ and $g$, the product $fg$ is defined by $f\circ g$ if the composition is possible and 0 otherwise. The ring (possibly without unity) $R[\mathcal{C}]$ has a natural structure of ring graded by the groupoid $Obj(\mathcal{C})\times Obj(\mathcal{C})$ via
@@ 2.4. Groupoid graded modules
A submodule $N$ of $M$ is a graded submodule if $N=\bigoplus_{\gamma\in\Gamma}N_{\gamma}$ where $N_{\gamma}:=N\cap M_{\gamma}$ for each $\gamma\in\Gamma$. Equivalently, if $n_{\gamma_{1}}+\cdots+n_{\gamma_{t}}\in N$ with $n_{\gamma_{i}}\in M_{\gamma_{i}}$ and $\gamma_{i}\neq\gamma_{j}$ for different i,j=1,,ti,j=1,...,t, then $n_{\gamma_{i}}\in N$ for all i=1,,ti=1,...,t.
@@
Let $M$, NN be $\Gamma$-graded right $R$-modules. A homomorphism of modules $g:M\to N$ it is said to be a gr-homomorphism of modules if $g(M_{\sigma})\subseteq N_{\sigma}$ for all $\sigma\in\Gamma$. We denote by $\operatorname{Hom_{gr-\mathit{R}}}(M,N)$ the abelian group consisting of the gr-homomorphisms of modules $g:M\to N$. We also define the ring $\operatorname{End_{gr-\mathit{R}}}(M):=\operatorname{Hom_{gr-\mathit{R}}}(M,M)$. It is easy to show that if $g\in\operatorname{Hom_{gr-\mathit{R}}}(M,N)$, then $\ker g$ is a graded submodule of $M$ and $\operatorname{im}g$ is a graded submodule of $N$. A bijective gr-homomorphism of modules $g:M\to N$ is called a gr-isomorphism of modules. We say that $M$ is gr-isomorphic to $N$ as modules (denoted by $M\cong_{gr}N$) if there exists a gr-isomorphism of modules $g:M\to N$.
@@
[11, Proposition 5] Let $R$ be a $\Gamma$-graded ring and let $M$ be a $\Gamma$-graded right $R$-module. Then $M$ is unital if and only if the equality $m_{\sigma}1_{d(\sigma)}=m_{\sigma}$ holds for all $\sigma\in\Gamma$ and $m_{\sigma}\in M_{\sigma}$. ∎
In this work, unless otherwise stated (Section 6), all groupoid graded modules are supposed to be unital. Moreover, in expressions such as “let $M$ be a $\Gamma$-graded $R$-module” we assume that $M$ is a $\Gamma$-graded right $R$-module.
@@
    Let $I$ be a set. Endow $I\times I$ with the groupoid structure of Example 2.1(2). Let $R$ be an $I\times I$-graded $R$-module. For each $i\in I$, recall that $1_{(i,i)}$ is the identity element of $R_{(i,i)}$. Fix $i_{0}\in I$. Then every unital right $R$-module $M$ can be regarded as an $I\times I$-graded $R$-module via $M_{(i_{0},i)}:=M1_{(i,i)}$ for each $i\in I$. Notice that $\operatorname{supp}(M)\subseteq\{i_{0}\}\times I$. Furthermore, any unital $(R,R)$-bimodule (i.e. $R\cdot M=M=M\cdot R$) can be regarded as an $I\times I$-graded $(R,R)$-bimodule via $M_{(i,j)}:=1_{(i,i)}M1_{(j,j)}$, for each $i,j\in I$. In particular this applies to the ring $R[\mathcal{C}]$ of a small preadditive category $\mathcal{C}$. In this case, every ideal of $R[\mathcal{C}]$ is $\mathcal{C}_{0}\times\mathcal{C}_{0}$-graded.
@@
Let $M$ be a $\Gamma$-graded right $R$-module and $\sigma\in\Gamma$. For each $\gamma\in\Gamma$, define $M(\sigma)_{\gamma}:=M_{\sigma\gamma}$, where we follow the convention $M_{\sigma\gamma}=\{0\}$ if $d(\sigma)\neq r(\gamma)$. The shift of $M$ by $\sigma$ is the $\Gamma$-graded right $R$-module $M(\sigma):=\bigoplus_{\gamma\in\Gamma}M(\sigma)_{\gamma}$. Note that $\operatorname{supp}(M(\sigma))\subseteq d(\sigma)\Gamma$. Analogously, if $M$ is a $\Gamma$-graded left $R$-module and $\gamma\in\Gamma$, set $((\sigma)M)_{\gamma}:=M_{\gamma\sigma}$. The shift of $M$ by $\sigma$ is the $\Gamma$-graded left $R$-module $(\sigma)M:=\bigoplus_{\gamma\in\Gamma}((\sigma)M)_{\gamma}$. Notice that $\operatorname{supp}((\sigma)M)\subseteq\Gamma r(\sigma)$. The following property of shifts of modules will be very useful. For a left version, see [11, Proposition 10(a)].
@@
Let $M$ be a $\Gamma$-graded $R$-module and $\sigma\in\Gamma$. Then $M(\sigma)$ equals $M(r(\sigma))$ as $R$-modules. More precisely, $M(\sigma)$ and $M(r(\sigma))$ have the same homogeneous components (albeit labeled with a different degree).
@@
Note that if $R$ is a $\Gamma$-graded ring and $M$ is a $\Gamma$-graded right $R$-module, then $M=\bigoplus_{e\in\Gamma_{0}}M(e)$ . This decomposition will be important, so we define
@@
The following results follow immediately from the definition (see [11, Proposition 13(c)]).